OpenStudy (ihats1):

mx+2nx=p solve for x

jhonyy9 (jhonyy9):

like a first step on the left side factorize out the x so what will get ?

OpenStudy (ihats1):

?? im at x+2nx=p-m is that correct? for the first step

OpenStudy (volleysappspbj16):

what he is saying is that mx and 2nx both share a factor of x so you can factor it out x(m+2n)=p

jhonyy9 (jhonyy9):

yes Volley this is the first step right - nice job

OpenStudy (volleysappspbj16):

:)

OpenStudy (ihats1):

so do i subtract x and divide p?

OpenStudy (volleysappspbj16):

no

jhonyy9 (jhonyy9):

and now after this for can separate the x just divide both sides by (m+2n)

OpenStudy (volleysappspbj16):

you can't subtract x from x(m+2n) because they are being multiplied, instead you must do the opposite of multiplication and divide like jhony9 said

OpenStudy (ihats1):

so divide both sides?

OpenStudy (volleysappspbj16):

yes

OpenStudy (volleysappspbj16):

by....

OpenStudy (ihats1):

no

OpenStudy (ihats1):

wait

OpenStudy (volleysappspbj16):

what do you divide both side by?

OpenStudy (ihats1):

x

OpenStudy (ihats1):

?

OpenStudy (volleysappspbj16):

no

OpenStudy (volleysappspbj16):

don't divide by x

OpenStudy (volleysappspbj16):

ever

OpenStudy (ihats1):

than what?

OpenStudy (volleysappspbj16):

or bad stuff will happen

OpenStudy (volleysappspbj16):

thats it

OpenStudy (volleysappspbj16):

x=p/(m+2n)

OpenStudy (ihats1):

x=p (m+2n)?

OpenStudy (volleysappspbj16):

no

OpenStudy (volleysappspbj16):

you divided both sides by (m+2n)

OpenStudy (ihats1):

so divide by p?

OpenStudy (volleysappspbj16):

no

OpenStudy (volleysappspbj16):

the answer is x= p divided by (m+2n)

OpenStudy (ihats1):

ok thank you

OpenStudy (volleysappspbj16):

uh hu

OpenStudy (triciaal):

here are the general guidelines first put all the terms with the subject on one side of the equation and move everything else to the other side. to move you do the inverse and whatever you do to one side of an equation you do to the other to keep the value. simplify each side if you can. now divide each side by the coefficient of the subject.

OpenStudy (triciaal):